### v2

# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

Abstract
In molecular aggregates, multiple delocalized exciton states interact with phonons, making the state-resolved spectroscopic monitoring of dynamics challenging. We propose a protocol that combines photon-entanglement-enhanced narrowband excitation of two-exciton states with time-frequency-filtered two-photon coincidence counting. This approach alleviates bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows. We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, thereby circumventing transport in mediating one-exciton states. The evolution of these population distributions and cascading transitions can be monitored using time-frequency-filtered photon coincidence counting. Numerical simulations for a light-harvesting aggregate highlight the ability of this protocol to achieve selectivity by suppressing or amplifying specific pathways. Combining entangled photonic sources with multidimensional photon counting allows promising applications in spectroscopy and sensing.
I Introduction
In interacting molecular aggregates, optical probing of dissipative exciton kinetics and the extraction of state-resolved spectroscopic information are inherently difficult. The high density of exciton states and exciton-phonon interactions lead to dissipative dynamics that span multiple time and energy scales. Molecular aggregates are found in naturally occurring photosynthetic systems. Within these photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport. Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic light-harvesting systems. [19, 1, 32]. In these systems, the aggregation of individual chromophores results in delocalized electronic excitations. These molecular excitons are primarily of the Frenkel type. The generated excitons participate in relaxation and dephasing, leading to primary charge-separation processes. Signature of exciton-exciton interactions and correlations, which are typically hidden in the dynamics of multi-exciton states were investigated [38, 35, 53]. The early attributes of such correlations can already be seen in the energy distributions of two-exciton states, which deviate from the sum of two one-exciton energies. Prior studies have demonstrated that various modalities of two-photon absorption spectroscopy can monitor the associated resonances. [54, 55, 23, 57, 29, 58, 8].
In recent years, several proposals restricted to atomic and simple molecular systems have underscored the advantages of employing entangled photon pairs to probe two-exciton states [51, 25, 49, 44, 18, 24, 9, 21, 22, 20, 48, 56, 34, 4, 6]. These proposals were complemented by critical examinations of the role of photon entanglement in enhancing two-photon absorption [47, 31, 30, 43, 46, 26, 2, 7]. Theoretical investigations have also explored the role of non-classical time-frequency correlations in obtaining superior resolution in condensed-phase spectroscopies. [14, 9].
Alongside, previous studies have proposed novel two-photon fluorescence intensity correlation measurements that can monitor the properties of two-exciton states [36, 33, 42, 15, 17, 50]. In a series of theoretical works, time-frequency filtered photon coincidence counting was proposed as a technique that may provide spectroscopic information similar to that of two-dimensional spectroscopy, but at different timescales [15, 17].
The optical excitation of two-exciton states in molecular aggregates, such as light-harvesting complexes, typically occurs via a manifold of one-exciton states subjected to population transport and dephasing. As a result, the population is redistributed among several states which may not be have highest propensity of transitions to the target two-exciton states. As the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well. Consequently, the number of interfering optical transitions increase dramatically. The final two-exciton population also undergoes transport, dispersing any initially created narrowband two-exciton distribution. In this work, we demonstrate that two-photon excitation using entangled photon pairs may achieve the creation of narrow-band population distribution. We also show that the time-frequency-filtered two-photon counting can characterize cascaded photon emission events, resulting two-dimensional correlation plots. It constitutes a spectroscopic technique that can serve as an efficient probe for dissipative exciton dynamics.
We focus on the light-harvesting complex LHCII, the most abundant antenna complex in plants and an important constituent of Photosystem II.
Figure 1: (A) Illustration of the proposed protocol, classified as a three-stage process: preparation of the two-exciton population, dissipative evolution of the two-exciton population, and time-frequency filtered detection of cascaded photon emission. The action of the time filter (governed by $F_{\bar{t}}^{(j)}$) precedes the action of the frequency filter (governed by $F_{\bar{\omega}}^{(j)}$) in both cases. The vertical straight arrows denote interaction with the driving photon modes, and the wiggly arrows denote interaction with the photon modes of the detector. In (B), three exemplery dynamical pathways are shown. The tilted double-ended arrows denote intra-manifold exciton transport. In (I), both two- and one-exciton transport preserves the narrowband nature of the created population distribution. In (II), during the excitation stage, the participation of one-exciton transport alters the target two-exciton population distribution. The population distribution disperses due to two-exciton transport, which results in broadband cascaded emissions. In (III), the narrowband two-exciton population distribution funnels to an even narrower distribution, but the subsequent one-exciton transport erases the memory. It lead leads to a narrowband emission followed by a broadband emission. The filtering functions assist in constructing a mechanistic narrative by classifying these pathways.
It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains several chlorophylls whose lowest three energy levels are spectroscopically accessible. In Section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII. We also the exciton-photon Hamiltonian. This is complemented by the Appendix A. In the subsequent sections, namely Sec. III (III.1III.3), we detail the theoretical framework for the combined protocol and numerical simulations. The manuscript concludes in Sec. IV with a summary of our findings and a brief commentary of the current limitations.
II Driven dissipative one- and two-exciton kinetics
We simulate the one- and two-exciton dynamics of the light-harvesting complex LHCII using the Frenkel exciton model within the Heitler-London approximation. The model is constructed by considering the lowest three electronic energy states for each chromophoric site. The Hamiltonian (with $\hbar=1$) is given by,
$H$ $=\sum_{mn}(E_{m}\delta_{mn}+J_{mn})B_{m}^{\dagger}B_{n}$
$+\sum_{mn}U_{m}^{(1)}B_{m}^{\dagger}B_{m}^{\dagger}B_{m}B_{m}+U_{mn}^{(2)}B_{m}^{\dagger}B_{n}^{\dagger}B_{n}B_{m}$
$+\sum_{j}\omega_{j}(b^{\dagger}_{j}b_{j}+\frac{1}{2})+\sum_{m,j}g_{m,j}(b^{\dagger}_{j}+b_{j})B_{m}^{{\dagger}}B_{m}$ (1)
where the exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$ (with $\eta=3/2$) and the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow the commutation relation $b_{i},b_{j}^{\dagger}]=\delta_{ij}$.
In the above, $E_{m}$ and $J_{mn}$ describes the on-site excitation and inter-site hopping, respectively. The later arises due the Coulomb interaction and leads to the delocalization of the exciton states across the sites. The one-exciton Hamiltonian can be constructed by considering the first term, i.e., $h_{mn}^{(1)}=(E_{m}\delta_{mn}+J_{mn})B_{m}^{\dagger}B_{n}$. Upon diagonalization, we obtain the corresponding one-exciton eigenstates $|e_{j}\rangle$.
The two-exciton Hamiltonian can be constructed by considering $h_{mnkl}^{(2)}=h_{mn}^{(1)}\delta_{kl}+\delta_{mn}h_{kl}^{(1)}+U_{m}^{(1)}\delta_{mk}\delta_{nl}\delta_{mn}+U_{mn}^{(2)}\delta_{mk}\delta_{nl}$ where $h_{mn}^{(1)}$ is given above. We note that the matrix elements of the two-exciton Hamiltonian differs from the sum of two one-exciton Hamiltonian by the last two terms. These terms which contribute to two-exciton nonlinearities, represent the energetic shift of overtone and combination exciton states respectively. In a similar manner as the previous case, the two-exciton eigenstates $|f_{k}\rangle$ are obtained by diagonalization.
The parameters for the Frenkel exciton Hamiltonian were taken from [41, 40, 10]. For the LHCII, the number of chromophoric sites is $N_{s}=14$, which yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states, composed of $14$ overtone and $91$ combination states. The eigenstates can be expressed as
$|e_{j}\rangle$ $=\sum_{m}T^{(1)}_{j,m}B_{m}^{\dagger}|0\rangle$
$|f_{k}\rangle$ $=\sum_{mn}T^{(2)}_{k,mn}B_{m}^{\dagger}B_{n}^{\dagger}|0\rangle$ (2)
where $T^{(1)}_{j,m}$ and $T^{(2)}_{k,mn}$ denote the matrix elements of the diagonalization matrices mentioned above. These matrix elements determine the spectral weights associated with the eigenstates which are delocalized over the site basis.
Figure 2: One- and two-exciton energies are displayed in the left and right panels, respectively. Nonuniform energy distributions are visible in both manifolds. Simulations in Section III.1 refer to the excitation of two-exciton states $f_{07}$ ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and $f_{83}$ ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) via $e_{07}$ ($E_{e_{07}}=15307.1\,\text{cm}^{-1}$) and $e_{09}$ ($E_{e_{09}}=15655.9\,\text{cm}^{-1}$).
The one exciton manifold spans from $E_{e_{1}}=14964.5\,\text{cm}^{-1}$ to $\omega_{e_{14}}=15934.7\,\text{cm}^{-1}$, hosting $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$. The two-exciton manifold spans from $E_{f_{1}}=29844.3\,\text{cm}^{-1}$ to $E_{f_{105}}=31811.8\,\text{cm}^{-1}$, hosting $105$ states within a spectral width of $\approx 1968\,\text{cm}^{-1}$. The corresponding energy distributions, plotted in Fig. 2, offers a visualization of the nature of the density of states in each manifold. We note that the energetic ordering of exciton eigenstates is dependent upon two competing parameters: $J_{mn}$ that govern exciton delocalization and $U_{m}^{(1)}$, $U_{mn}^{(2)}$ that govern exciton nonlinearities.
The dissipative exciton dynamics is described by the Liouville-space Green’s function formalism. Using the exciton eigenbasis, we can express the Green’s function as follows
$\mathcal{G}_{a_{1}a_{2},a_{3}a_{4}}(t)=\delta_{a_{1}a_{2}}\delta_{a_{3}a_{4}}\theta(t)[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
$+(1-\delta_{a_{1}a_{2}})\delta_{a_{1}a_{3}}\delta_{a_{2}a_{4}}\theta(t)\exp{(-i\omega_{a_{1}a_{3}}t-\gamma_{a_{1}a_{3}}t)}$ (3)
The expression above is valid for both exciton transport and dephasing, which are captured by the first and the second terms, respectively. It is discussed further in Appendix A.
Interaction with the photonic sources, within the rotating-wave approximation, can be described by the following Hamiltonian:
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{\Omega}}$
$\Big(a_{j}\exp{(-i\omega_{j}t)}\exp{(i\omega_{pq}t)}B_{pq}+\mathrm{h.c.}\Big)$ (4)
In the above, we used a mode expansion of the photonic field operators, where $a_{j}$ denotes the photon annihilation operators and $\omega_{j}$ represents the mode frequencies. These mode operators follow bosonic commutation relations. The exciton annihilation operators, expressed in the eigenbasis, are denoted as $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$. The entangled photon pairs induce inter-manifold exciton transitions.
Figure 3: (A) The Schwinger-Keldysh loop diagram describing the pathways involved in the proposed protocol. The excitation and emission states involve four exciton-photon interactions each with the driving and detection photon modes. The Liouville space ladder diagrams in (B)-(C) are suitable for capturing the processes in dissipative systems. The excitation and emission pathways are depicted in (B) and (C) respectively. In the excitation stage, described in (C), diagrams (II) and (IV) depends on one-exciton transport, $|e\rangle\langle e|\rightarrow|e^{\prime}\rangle\langle e^{\prime}|$. Diagram (I) depends on inter-manifold two-exciton coherence, $|f\rangle\langle g|$, while (III) and (V) depend on intra-manifold one-exciton coherence, $|e\rangle\langle e^{\prime\prime}|$. In the emission stage, described in (B), all diagrams have the two-exciton population distribution following transport as the initial state, $|f\rangle\langle f|\rightarrow|f^{\prime}\rangle\langle f^{\prime}|$. A combinatorial pairing of five excitation and four possible emission diagrams generate all possible pathways that may contribute to the signal. Their relative contributions to the final signal is jointly determined by the four-point photon correlation functions and the filtering properties. The corresponding equations are presented in Section III.
We consider entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit. In a typical scenario, a birefringent crystal is pumped with a classical laser, leading to the generation of correlated photon pairs with frequencies $\omega_{1}$ and $\omega_{2}$. The SPDC process involves two photon-matter interactions with the incoming classical field modes and four photon-matter interactions with the outgoing signal-idler fields. The correlation properties of the emitted photon field are determined by the physical properties of the SPDC crystal. [49, 52, 25, 51, 16, 9, 10].
The generated entangled two-photon states can be described as follows:
$\rho_{\text{photon}}$ $=\int\frac{d\omega_{1}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}^{\prime}}{2\pi}\int\frac{d\omega_{2}^{\prime}}{2\pi}$
$F(\omega_{1},\omega_{2})F^{*}(\omega_{1}^{\prime},\omega_{2}^{\prime})a_{\omega_{1}}^{{\dagger}}a_{\omega_{2}}^{{\dagger}}|0\rangle\langle 0|a_{\omega_{1}^{\prime}}a_{\omega_{2}^{\prime}}$ (5)
The four-point correlation function of the entangled photon pairs is relevant for the creation of two-exciton states. Using the expression above, it can be evaluated, yielding
$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle_{\rho_{\text{photon}}}$
$=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
We defined, $F(\omega_{a},\omega_{b})=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$, where $\phi(\omega_{a},\omega_{b})=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ and $\alpha$ denotes the parameter governing the down-conversion efficiency. The entanglement time parameter $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establishes an upper bound for the intervals between two photon-exciton interactions and depends on the time delays acquired by the signal and idler photons during propagation within the nonlinear crystal. It characterizes the non-classical correlation in the photonic source. The classical transform-limited pulse driving the SPDC crystal is defined as follows: $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width.
III Results
The sequence of processes: entangled photon excitation, dissipative evolution, and two-photon emission occurs in three stages, as depicted in Fig. 1 (A) (I) to (III). Excitation by entangled photon pairs creates a tailored two-exciton population distribution which undergoes transport for a finite time. It results in a redistribution of populations. The redistributed population serves as the initial configuration for the cascaded two-photon emission processes. The emission process involves two successive one-photon emission events interleaved by one-exciton transport in the one-exciton manifold.
In Fig. 1 (B), we demonstrate three plausible pathways via which the dynamics may proceed. Due to transport, the two-exciton population distribution may lose its narrowband character or may even become narrower. The dynamical information is encoded into the photon emissions that correspond to the inter-manifold transitions between the upper two manifolds ($|f\rangle\rightarrow|e\rangle$). The one-exciton transport that follows may also create additional dispersion of the one-exciton population. The corresponding dynamical information is encoded into the inter-manifold transition between lower two manifolds ($|e\rangle\rightarrow|g\rangle$). These two sets of photon emissions can be classified by controlling the properties of the time and frequency filters, denoted $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$ in the Fig. 1.
In the subsequent sections, we use the following notation for Liouville-space operators: $A_{\alpha}$, where the subscript denotes $\alpha=L,R,+,$-. These operators are defined by their action on a Hilbert-space operator $X$ as $A_{L}X\equiv AX$ and $A_{R}X\equiv XA$. The matrix elements are obtained as $(A_{\alpha}X)_{ij}\equiv\sum_{kl}(A_{\alpha})_{ij,kl}X_{kl}$. We also define $A_{+}\equiv\frac{1}{2}(A_{L}+A_{R})$ and $A_{-}\equiv A_{L}-A_{R}$, where the matrix elements are given by: $(A_{-})_{ij,kl}=A_{ik}\delta_{jl}-A_{\ell j}\delta_{ik}$ and $(A_{+})_{ij,kl}=\frac{1}{2}[A_{ik}\delta_{jl}+A_{lj}\delta_{ik}]$.
The time-frequency filtered two-photon coincidence signal can be expressed as follows:
$S(\overline{t}_{s},\overline{\omega}_{s};\overline{t}_{i},\overline{\omega}_{i})$ $=\int_{-\infty}^{\infty}dt_{s}^{\prime}d\tau_{s}\int_{-\infty}^{\infty}dt_{i}^{\prime}d\tau_{i}\sum_{s,s^{\prime},i,i^{\prime}}\text{Tr}_{s,s^{\prime},i,i^{\prime}}[D^{(s,s^{\prime})}(\overline{t}_{s},\overline{\omega}_{s};t_{s}^{\prime},\tau_{s})D^{(i,i^{\prime})}(\overline{t}_{i},\overline{\omega}_{i};t_{i}^{\prime},\tau_{i})B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})]$ (7)
Here we define the superoperator for filtered detection:
$D^{(\nu,\nu^{\prime})}(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})$
$=D(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})E_{\nu^{\prime}L}(t_{\nu}^{\prime})E_{\nu R}^{\dagger}(t_{\nu}^{\prime}+\tau_{\nu})$ (8)
and for the bare two-photon signal
$B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})$
$=T_{L}^{(s^{\prime},i^{\prime})}(t_{s}^{\prime},t_{i}^{\prime})T_{R}^{(s,i)\dagger}(t_{s}^{\prime}+\tau_{s},t_{i}^{\prime}+\tau_{i})$ (9)
We also use the following superoperators to define the emission processes:
$T_{R}^{{(s^{\prime},i^{\prime})},{\dagger}}(t_{s}^{\prime},t_{i}^{\prime})$ $=i^{2}\int_{-\infty}^{t_{2}^{\prime}+\tau_{2}}dt_{4}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
$E_{sR}(t_{3})E_{iR}(t_{4})V_{R}^{{\dagger}}(t_{4})V_{R}^{{\dagger}}(t_{3})$
$T_{L}^{(s^{\prime},i^{\prime})}(t_{s}^{\prime},t_{i}^{\prime})$ $=i^{2}\int_{-\infty}^{t_{2}^{\prime}}dt_{2}\int_{-\infty}^{t_{1}^{\prime}}dt_{1}$
$E_{sL}^{\dagger}(t_{1})E_{iL}^{\dagger}(t_{2})V_{L}(t_{1})V_{L}(t_{2})$ (10)
Here the superoperators $T_{L}$ and $T_{R}$ are defined by their action from the left and right sides of the density operator, which describes the initial configuration for the photon emission following the excitation and two-exciton transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$. For future convenience, we also define the observable that corresponds to the preparation of the two-exciton population distribution
$O(t)=\tilde{T}_{R}^{{\dagger}}(\tau_{2},\tau_{3})\tilde{T}_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
$\int_{-\infty}^{\tau_{3}}d\tau_{2}\int_{-\infty}^{\tau_{4}}d\tau_{1}E_{L}^{\dagger}(\tau_{4})E_{L}^{\dagger}(\tau_{1})E_{R}(t_{3})E_{R}(\tau_{2})$
$\qquad\qquad V_{R}^{{\dagger}}(\tau_{3})V_{R}^{{\dagger}}(\tau_{2})V_{L}(\tau_{4})V_{L}(\tau_{1})$ (11)
Going forward, we assume that the excitation and emission stages are temporally non-overlapping, which results in the factorization of Eq. 7 into separate contributions from each of the three stages. In the following sections, we systematically analyze the role of each stage.
Figure 4: The two-exciton population distribution map, constructed by sequentially targeting each of the $105$ states in that manifold, using degenerate entangled photon pairs i.e., $\omega_{i}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. The left panel serves as a reference simulation while the temporal width of the SPDC pump (entanglement time) is varied in the middle panel (right panel). Dispersal of the plot, in the horizontal cut, away from diagonal signifies a reduction in the narrowband excitation. For discussion, see section III.1
III.1 Entangled photon-pair excitation
The quantity of interest at the preparation stage, the two-exciton population distribution, is defined by the projection operator $|f\rangle\langle f|$ acting on the two-exciton density operator, expressed as $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]$, where the partial trace $\text{Tr}_{B}$ is taken over the phonon modes. We perturbatively expand the density operator to the fourth order in $H_{\text{int}}(t)$ to obtain the expression for the observable:
$\rho_{ff}(t)$ $\propto 2\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}^{*}d_{eg}^{*}\times I(-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1};-z_{ff})I(-\omega_{3}+\omega_{2}+\omega_{1};-z_{fe^{\prime}})I(+\omega_{2}+\omega_{1};-z_{fg})I(\omega_{1};-z_{eg})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1};-z_{ff})I(\omega_{2}-\omega_{3}+\omega_{1};-z_{fe^{{}^{\prime\prime}}})$
$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1};-z_{ff})I(-\omega_{4}-\omega_{3}+\omega_{1};-z_{fe^{\prime\prime}})$
$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (12)
The derivation is outlined in Appendix C. The expression above is a sum of five distinct terms, each representing an exciton pathway that contributes to the signal. The Liouville-space Feynman diagrams in Fig. 3 (I)–(V), offer further insights into the exciton processes captured by this expression. The pathways include contributions from both exciton coherences and populations. The first term represents the two-exciton coherence-mediated pathway (diagram (I)), which is susceptible to inter-manifold dephasing. The second and fourth terms describe pathways susceptible to one-exciton population transport (diagrams (II) and (IV)). The third and fifth terms describe pathways susceptible to intra-manifold dephasing (diagrams (III) and (V)).
The convolutional form provided above is particularly advantageous for exploring the role of entangled photon pairs. Their effect is included via the four-point correlation functions, defined in Eq. II. The following two sections use both degenerate and nondegenerate entangled photon pairs to illustrate these capabilities.
III.1.1 Simulation: excitation of the full suite of two-exciton states using degenerate entangled photon pairs
In this section, we monitor the final two-exciton population distribution following excitation by spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. For an SPDC source pumped by a finite-bandwidth laser, the generated photon pairs usually carry a range of frequencies. The photon pairs, emitted within a specific solid angle, must be filtered to select the degenerate components. We carry out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig. 4. The results presented in the left column use an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, which acts as a reference case. The middle and right columns present simulations with a variation in the temporal width of the SPDC pump ($\tau_{0}=50\,\mathrm{fs}$) and the entanglement time parameter ($T_{\mathrm{ent}}=60\,\mathrm{fs}$), respectively, while keeping all other parameters unchanged.
We note that in all three cases, when the target states are located near the middle of the two-exciton manifold, the results exhibit broader population distributions around the target. In other words, we find a lower degree of selectivity. In comparison, when the target states are located in the lower-energy sector of the two-exciton manifold, we observe a higher degree of selectivity. Altering the temporal width of the SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways. In contrast, altering the entanglement time in the right column preserved the selectivity for the majority of states. Across all plots, we notice a trend: some two-exciton states remain preferentially populated across all parameter regimes. This feature has likely emerged from an intricate interplay of exciton transport parameters of mediating states.
Figure 5: The population distribution for the target state $f_{07}$ which is dominantly excited via the intermediate states $e_{07}$ and $e_{09}$ is displayed for variations in the entangled photon parameters (rows) and for two distinct transport times (columns): $t=0\,\text{fs}$ (immediately after the pulse) and $t=50\,\text{fs}$. In the upper panel, the entanglement time is $\tilde{T}_{\text{ent}}=10\,\text{fs}$ and the SPDC pump width is $\tau_{0}=150\,\text{fs}$. In the middle panel (lower panel), the width is reduced (entanglement time is increased) to $\tau_{0}=50\,\text{fs}$ ($\tilde{T}_{\text{ent}}=30\,\text{fs}$). For detailed discussion, refer to Section III.1.
III.1.2 Simulations: differential preparation of two-exciton populations via preselected mediating states and parametric dependence on photonic properties
We focus on two specific two-exciton states among the full suite of target states previously monitored, namely, $f_{07}$ and $f_{83}$. These two-exciton states reside in the lower ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and the higher ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) energy regions of the two-exciton manifold, respectively. We examine the possibility of exciting these target states via a fixed pair of one-exciton states, selected by setting $\omega_{1}=E_{e_{07}}=15307.1\,\text{cm}^{-1}$ and $\omega_{2}=E_{e_{09}}=15655.9\,\text{cm}^{-1}$. These mediating states are delocalized over multiple chromophoric sites and are prone to ultrafast transport.
These results are displayed in the left columns of Fig. 5 and Fig. 6, respectively. The entanglement time and temporal width of the SPDC pump are set to $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$ for the upper panel, which is similar to the reference case in Fig. 4. In the middle panel, the temporal width is changed to $\tau_{0}=50\,\mathrm{fs}$, whereas in the bottom panel, the entanglement time is changed to $T_{\mathrm{ent}}=40\,\mathrm{fs}$.
The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy in some cases. The increase in the temporal width results in a significant loss of selectivity as the number of mediating pathways with comparable spectral weights grows. The variation of the entanglement time didn’t result in any major changes, except for the tailoring of the relative populations. The latter, however, will have consequences as the population redistributes over time, as will be shown in the next section. We emphasize that both the target states are excited via the same mediating stating one-exciton states. Therefore the differential nature of the final population distribution is dependent on the transition dipoles that connects chosen one-exciton states to the two-exciton target states. The final population distribution demonstrates that the short entanglement time reaches a saturation, less so for the $f_{07}$ than $f_{83}$. The higher lying transitions for the latter imposes stringent requirements on the frequencies of the entangled photon pairs.
III.2 Transport in the two-exciton manifold
The two-exciton population distribution evolves in time, governed by the properties of the transport matrix. The following questions merit further investigation: How does the initial two-exciton population distribution redistribute itself over time? To what extent the two different population distributions differ after undergoing redistribution? Finally, to what extent does the redistribution dynamics depend on the properties of the entangled photon source?
The kinetic equation that governs the dissipative population dynamics of the two-exciton populations is given by
$\frac{d}{dt}\rho_{ff}(t)=-\sum_{f^{\prime}}K_{ff,f^{\prime}f^{\prime}}\rho_{f^{\prime}f^{\prime}}(t)$ (13)
where the transport matrix, $K_{ff,f^{\prime}f^{\prime}}$, follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$. The solution to this equation yields the transport Green’s function $\mathcal{G}_{ff,f^{\prime}f^{\prime}}(\tau)$. The initial two-exciton population distribution, obtained via Eq. III.1, is propagated using the following expression: $\rho_{ff}(t_{2})=\mathcal{G}_{ff,f^{\prime}f^{\prime}}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$. The transport Green’s function is given by the first term in Eq. II and is discussed further in Appendix A.
The following four sets of simulations will help us address the questions posed above.
III.2.1 Simulations: differential evolution of two-exciton populations and parametric dependence on photonic properties
We monitor the nature of differential population evolution after the excitation targeted at $f_{07}$ (right column in Fig. 5) and $f_{83}$ (right column in Fig. 6), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively. Each of the three panels, upper, middle and bottom, represents simulation using different photonic parameters. The parameters were listed in the preceding section III.1.2 where we discussed the reference simulation. We observe that the evolved distributions at short times retained the memory of prepared distribution in Fig. 5, but gets washed out in Fig. 6. The effect of the variation of photonic parameters is also less visible in the latter. The rapid redistribution and localization towards a few selected states in the latter is a key feature that requires future investigation. It can also be concluded that such a prepared distribution would loose the principal advantages of the entangled photon pair excitations.
Figure 6: The population distribution for the target state $f_{83}$ which is dominantly excited via the intermediate states $e_{07}$ and $e_{09}$ is displayed for variations in the entangled photon parameters (rows) and for two distinct transport times (columns): $t=0\,\text{fs}$ (immediately after the pulse) and $t=250\,\text{fs}$. In the upper panel, the entanglement time is $\tilde{T}_{\text{ent}}=10\,\text{fs}$ and the SPDC pump width is $\tau_{0}=150\,\text{fs}$. In the middle panel (lower panel), the width is reduced (entanglement time is increased) to $\tau_{0}=50\,\text{fs}$ ($\tilde{T}_{\text{ent}}=30\,\text{fs}$). For detailed discussion, refer to Section III.1.
III.2.2 Simulations: differential redistribution of two-exciton populations
In Fig. 7, we track the population redistribution by displaying the population snapshots at $t=50,100,250,1000\,\text{fs}$ following the excitation targeted at $f_{07}$ (in the left column) and $f_{83}$ (in the right column). The results demonstrate two distinct features: the population migrates to the lower energy sector of the manifold rather monotonically for $f_{07}$ and by a slower intra-manifold cascading mechanism for $f_{83}$. The slower transport can be attributed to lack of overlap between the latter with the state in the middle rung of two-exciton manifold.
The population redistribution can be spectroscopically monitored by combining spectral-temporal filtering schemes in the detection of the two-photon emission. [16, 14, 39]. Such a scheme will track cascaded inter-manifold transitions fluorescence that competes with the intra-manifold transport. The next section is devoted to the theory and simulations demonstrates these schemes.
Figure 7: Population redistribution at four different time points following the excitation of the target states $f_{07}$ (right column) and $f_{83}$ (left column). The results demonstrate key features of the dynamics, including rapid funneling of the population along energy gradients and cascaded population migration for the two targets, respectively. For detailed discussion, refer to Section III.2
III.3 Photon correlation spectroscopy
The quantity of interest in photon correlation spectroscopy is the time–frequency filtered two-photon counting signal. In this section, we describe how it can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways.
The rationale for deploying a filtering scheme can be developed as follows. Population transport in the two-exciton manifold acts to redistribute the initially created population. The timescale of this phenomenon is determined by the duration of transport. The population redistribution has two principal consequences: it gives rise to new set of inter-manifold pathways and reduces the spectral intensity of some transitions. Each photon emission pathway can be associated with a specific time and frequency window, which can be selectively monitored by tuning the properties of the first set of time and frequency filters. Following photon emission, the dynamics continue in the one-exciton manifold, where transport may scramble the population distribution. After a finite duration, another inter-manifold transition leads to the ground state. Similar to the above, these transitions can also be monitored via appropriate filtering. Three pathways in Fig. 1 (B) show the dynamical complexities accompany each photon emission. In systems with exciton manifolds, multiple photon emission pathways are possible.
The signal is expressed as:
$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
$\times\int dt_{2}^{\prime}\int d\tau_{2}D^{(2)}(t_{2},\omega_{2};t_{2}^{\prime},\tau_{2})\langle N_{r}(t_{1}^{\prime},\tau_{1})N_{s}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle$
$=\int dt^{\prime}_{1}\int d\tau_{1}\int dt^{\prime}_{2}\int d\tau_{2}D^{(1)}(t_{1}^{\prime},\tau_{1})D^{(2)}(t_{2}^{\prime},\tau_{2})$
$\sum_{s,s^{\prime}}\sum_{i,i^{\prime}}\langle E_{s^{\prime}}^{\dagger}(t^{\prime}_{2}+\tau_{2})E_{i^{\prime}}^{\dagger}(t^{\prime}_{1}+\tau_{1})E_{i}(t^{\prime}_{1})E_{s}(t^{\prime}_{2})\rangle$ (14)
Here, $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}E_{sR}^{\dagger}(t^{\prime}+\tau)E_{s^{\prime}L}(t^{\prime})$ is a photon number operator without the filter. The field operators can be expanded as: $E_{s}(t)=\sqrt{2\pi\omega_{s}/\Omega_{s}}\hat{a}_{s}e^{-i\omega_{s}t}$, where $\Omega$ is the mode quantization volume for the photonic modes of the detector. The latter follows the Bosonic commutation relation $a_{s},a_{s^{\prime}}^{\dagger}]=\delta_{ss^{\prime}}$. We define two spectrogram functions, $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$, which are governed by the time and frequency filtering functions, $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$. The spectrogram function can be represented as
$D(\bar{t},\bar{\omega},t^{\prime},\tau)$
$=\int\frac{d\omega^{\prime\prime}}{2\pi}e^{-i\omega^{\prime\prime}\tau}|F_{\bar{\omega}}(\omega^{\prime\prime},\omega)|^{2}F_{\bar{t}}^{*}(t^{\prime}+\tau,t)F_{\bar{t}}(t^{\prime},t).$ (15)
We choose the following time and frequency filters: $F_{\bar{t}}(t^{\prime})=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ and $F_{\bar{\omega}}(\omega^{\prime})=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$. The spectrogram function can be evaluated as
$D(\bar{t},\bar{\omega},t^{\prime},\tau)=\frac{1}{2\sigma_{\omega}}\theta(t^{\prime}-t)\theta(t^{\prime}+\tau-t)$
$[\theta(\tau)e^{-\sigma_{\bar{\omega}}\tau}+\theta(-\tau)e^{\sigma_{\bar{\omega}}\tau}]e^{-(i\omega+\sigma_{\bar{t}})\tau-2\sigma_{\bar{t}}(t^{\prime}-t)}$ (16)
It depends on four parameters: the centering times and frequencies of the time and frequency filters, denoted $\bar{t}_{j}$ and $\bar{\omega}_{j}$, and their respective widths, $\sigma_{\bar{t}_{j}}$ and $\sigma_{\bar{\omega}_{j}}$. One can modulate the spectrogram by varying these parameters. The spectral filters resolve inter-manifold transitions that originate from exciton transport occurring within similar time windows. The temporal filters help separate the transport processes that may involve similar spectral gaps. Combined, they monitor the inter-manifold transitions jointly in time and frequency space.
Using the expression of the spectrogram and following the steps described in Appendix E, we can express the signal as a sum of four pathways (which is reduced to two following symmetry arguments)
$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\sum_{p=1}^{4}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})$
$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
$\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime}-\tau_{2})\mathcal{G}_{ef^{\prime},ef^{\prime}}(\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime})|\rho_{ff}\rangle\rangle$
$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
$\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime})\mathcal{G}_{f^{\prime}e,f^{\prime}e}(-\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime}+\tau_{2})|\rho_{ff}\rangle\rangle$ (17)
In the above, subscripts in the expression for the spectrograms, $D_{>}$ and $D_{<}$, indicate the sign of the delay $\tau$ between two emissions. The density of photonic states of the detector is assumed to be concentrated at frequencies corresponding to the centering frequency of the frequency filter; these are denoted as $\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})$. In Fig. 3 (I)–(IV), we diagrammatically illustrate the emission pathways in Liouville space. We note that each diagram corresponds to pathways that involve transport in both manifolds, duration of which is controlled by the centering times of the time filters.
Figure 8: The time-frequency filtered two-photon counting signal is displayed. The initial initial population distribution for the photon counting is given by the targeted excitation of $f_{07}$ followed by two-exciton transport. Two sets of time and frequency filters can allow monitoring of the subsequent processes such as inter-manifold transition, one-exciton transport, and inter-manifold transition. The reference simulation in the upper-left uses the following parameters: width of the frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, width of the time filters $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$. In upper-right the the width of the frequency filter is varied to $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$. In middle-left the centering time of the time filter is varied to capture $t_{w,2}=1000\text{fs}$, whereas in middle-right we vary widths of frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ while keeping the and the centering time of the time filter same as middle-left; In bottom-left, the width of the time-filter is varied to $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$. In bottom-right, the centering time of the time filter is varied to capture $t_{w,1}=50\text{fs}$. For detailed discussion, see Section III.3.1
III.3.1 Simulation: two-photon coincidence counting signal with variation in filtering parameter and different waiting times
In Fig. 8, we present six two-dimensional plots where centering times of the frequency filters (i.e., $\bar{\omega}_{j}$) are taken as the scanning parameters. These results explore the signal for cases where the widths (i.e., $\sigma_{\bar{t}_{j}}$) and centering parameters of time filters (i.e., $\bar{t}_{j}$) are varied. We note that, the latter help unravel the exciton transport timescales. The two-exciton distribution targeted at $f_{07}$, after a period of time evolution serves as the initial configuration. The peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two-exciton to and one-exciton manifold $|f\rangle\rightarrow|e\rangle$, while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state $|e\rangle\rightarrow|g\rangle$. The former monitors the spectral gaps $\sim\omega_{fe}$ while the latter the spectral gaps $\sim\omega_{eg}$. This can be ensured by choosing the centering times of the time filter $\bar{t}_{2}$ well-ahead of $\bar{t}_{1}$.
As highlighted earlier in Section II, density of energy gaps is significantly higher in the upper set of transitions. We recall that the one-exciton manifold hosts $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$ and following excitation of $f_{07}$, the two-exciton manifold hosts $\sim 10$ states within a spectral width of $\approx 350\,\text{cm}^{-1}$. The number of plausible transitions remains large $\sim 140$. Two filtering stages allow classify these transitions in different temporal windows.
For the upper-left panel, the reference simulation for this set, we use the following parameters: $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$. Variables $t_{\text{w},j}$ are associated with the $\tau_{j}$ integrals in the expression of detector spectrograms. The waiting times determine the duration for which exciton transport is allowed at each manifold, before the time-filter is opened. In the plot, we observe eight resonance zones spread along the $\bar{\omega}_{1}$ axis, each containing several resonances that can be further resolved along the $\bar{\omega}_{2}$ axis. For easier interpretation, we may classify the resonances along $\bar{\omega}_{1}/10^{3}\,\text{cm}^{-1}$ in four distinct clusters, each containing two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$) resonances. Since only the transition energy differences are being captured, each of these transitions, in the absence of the time filter would have appeared in a congested form. The four observed clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively. For each of these clusters, the resonances along $\bar{\omega}_{2}/10^{3}\,\text{cm}^{-1}$ are identified as $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively.
The upper-right panel repeats the simulation with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the previous one. We observe, along the $\bar{\omega}_{1}$ axis, that several resonance zones in the second cluster are absent; the most noticeable absences occur in the middle and low energy sectors. Additionally, a few resonances along the $\bar{\omega}_{2}$ are absent. The fourth cluster has been diminished in intensity. The broadband spectral filters used in the simulation increases the number of destructive interferences. Given that the transitions originate from a limited number of states number of states it also highlights that they are preferentially transitioning to a few selected one-exciton states.
The middle-left panel investigates the effect of longer waiting times aimed at identifying the role of one-exciton transport. We set $t_{\text{w},2}=1000\,\text{fs}$ while keeping all other parameters unchanged from the reference simulation. The population distribution in the one-exciton manifold depends on two factors: the dipole-connectivity of states with those of the two-exciton manifold and the transport rates. We note that along the $\bar{\omega}_{1}$ axis the clusters have further coalesced. They are clearly separated into two dominant regions. Furthermore, the fourth cluster has vanished, and the resolution along the $\bar{\omega}_{2}$ axis is lost. The observation indicates that, at longer waiting times, the connectivity dictated by transition dipoles focuses the spectral weights to a few selected states.
The middle-right panel repeats the reference simulation but with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$. Therefore it is aimed at exploring the role of a broadband frequency filter on one-exciton transport at longer times. We note the resonances along the $\bar{\omega}_{2}$ axis and conclude that widening the spectral gate may not have permitted a larger number of transitions. The reasons are twofold: the initial two-exciton distributions, being restricted to a limited set of states, have no further transitions to explore and these transitions at longer times participates in dynamics that converges to same set of states.
The bottom-left panel explores the examines the role of the width of the time filter by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ while keeping all other parameters the same as the reference. We note that the second cluster has a diminished presence. Also, several resonances along the $\bar{\omega}_{2}$ axis are absent. We note that positioning the centering time of the filter at shorter time effectively skips the resonances that are generated as a result of the longer time dynamics. Hence, by comparing the reference simulation and this one, one may identify the transitions that takes place at longer times.
The bottom-right panel examines the role of the two-exciton transport by setting $t_{\text{w},1}=50\,\text{fs}$. All other parameters are kept as the same as that of the reference. We observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structure, while there is no significant change along the $\bar{\omega}_{1}$ axis. The longer two-exciton transport tend to funnel all populations to a few low-lying two-exciton states, as can be observed in Fig. 5 and Fig. 7. This makes a fewer number of transitions possible, reducing the number of corresponding resonances. We note that the for a robust phenomenological description the state indices from the figures need to be converted to energy values.
IV Conclusions
In this work, we demonstrated that a combining entangled photon-pair excitation and time-frequency filtered two-photon counting provides a systematic route to probe dissipative one- and two-exciton kinetics. The combined protocol was applied to the interacting molecular aggregate, light-harvesting complex II (LHCII), where exciton–exciton and exciton–phonon interactions jointly govern to the dynamics. The protocol offers a classification scheme for exciton dynamics, even when they span multiple time and frequency scales.
Results in the Section III.1 reveals that a narrow-band population distribution around some of the two-exciton states can be created with a high degree of fidelity. The final population distribution is examined against variations in the entangled photon parameters and the temporal width of the SPDC pump. We also examined the differential nature of excitation by targeting two distinct two-exciton states. Because these two target states are excited via the same one-exciton states, they offer insight into pathway selectivity. From the current simulations, a strategy emerges: preferentially accessing the one-exciton resonances that are less susceptible to transport is the key to achieving robust, state-selective excitation of two-exciton states. The expression in Eq. III.1 remains valid for a wide range of photonic sources, provided the dynamics remains within the perturbative regime of exciton–photon coupling. The expression consists of five terms, each containing a convolution of exciton Green’s functions with four-point correlation functions of the driving photonic sources. They modulate spectral weights for each pathway. Our future work will address how multiobjective optimization algorithms can be used to actively tailor these correlation functions.
The evolution of the initially created two-exciton population is studied in Section III.2. This evolved population distribution serves as the initial configuration for the photon emission pathways. The presence of transport within the two-exciton manifold effectively scrambles the identity of the states initially created by two-photon absorption. Similarly, transport in the one-exciton manifold have the same effects on the state distribution generated after the first photon emission. Interestingly, the one-exciton transport may exert the scrambling effect on two occasions: during the excitation and again during the emission process.
In Section III.3, we deployed the time-frequency filtering to track inter-manifold transitions. The expressions in Eq. III.3 remain valid for general two-exciton population distributions. The variation of the filtering parameters may offer detailed insights into spectral gaps, intra-manifold transport timescales, and inter-manifold dephasing. We assumed that the excitation and emission events are temporally separated which omits spontaneous Raman contributions. The joint spectral-temporal filtering effectively enhances the resolution of the bare two-photon counting signal. Two spectrograms, parametrized by filtering functions separate the exciton pathways jointly in time and frequency. It allows one to construct a kinetic description of state-resolved exciton dynamics even in the presence of dissipation.
The entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of successive photon-exciton interactions. In contrast, the bandwidth of the classical spontaneous parametric down-conversion (SPDC) pump determines the range of accessible resonances. The independent controllability of these variables may affords additional flexibility. For a given $N$-site exciton aggregate, there are $N$ one-exciton states and N+N(N1)/2N+N(N-1)/2 two-exciton states. As the size of the aggregate increases, the dynamics may span a broad range of time and frequency values. To probe such dynamics a wider tunability of photonic parameters may become necessary [5, 37, 11]. Similar constraints apply on the filtering parameters, which must be able to characterize a wide range of two-photon emission pathways.
In general, entangled photon pairs offer two primary advantages for two-photon spectroscopy: the ability to modulate competing excitation (and de-excitation) pathways to selectively suppress or amplify specific exciton resonances, and favorable scaling of the signal with the intensity of the incoming field. We note that the correlation function of coherent laser fields factorizes into a product of four amplitude functions, i.e., $\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{2}(\omega_{2})A_{1}(\omega_{1})$. The signal scales with the square of the incident field intensity. In contrast, for the entangled photon pairs, the signal scales linearly. However, realizing these benefits remains a significant challenge due to the inherently low signal-to-noise ratio and the difficulty of separating contributions arising from photonic entanglement [45, 26, 27]. Technological advances in this direction would establish the proposed protocol as a suitable tool for investigating non-classical effects in light-harvesting complexes.
The time- and frequency-filtered $n$-photon counting signal can analyze multiphoton emission in a $2n$-dimensional parameter space. If combined with an efficient state preparation using multiphoton entangled sources, it may provide a robust spectroscopic tool for characterizing dissipative molecular excitonic systems. Such approaches can also be extended to programmable photonic sources and interferometric detection schemes [28, 13, 3].
Acknowledgements.
A.D. acknowledges the support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF. S.M. gratefully acknowledges the support of the National Science Foundation (NSF) grant Grant No. CHE-2246379.
Appendix A The Hamiltonian
In the Hamiltonian, the Eq. II, the vibrational terms are specified in the third line. The normal modes associated with collective intermolecular vibrations constitute the phonon modes. The latter, in the continuum limit, i.e., $J_{\mathrm{ph}}(\omega)=\sum_{j}|g_{j}|^{2}[\delta(\omega-\omega_{j})-\delta(\omega+\omega_{j})]$ are described using an overdamped oscillator and $48$ multimode Brownian oscillators. The corresponding spectral function is given by
$J_{\mathrm{ph}}(\omega)$ $=2\lambda_{0}\frac{\omega\gamma_{0}}{\omega^{2}+\gamma_{0}^{2}}+\frac{2\lambda_{j}\omega_{j}^{2}\omega\gamma_{j}}{(\omega_{j}^{2}-\omega^{2})^{2}+\omega^{2}\gamma_{j}^{2}}$ (18)
Here, $\lambda_{0}$ and $\gamma_{0}$ represent the damping strength and relaxation parameter for the overdamped mode, respectively. The Brownian oscillators are characterized by the Huang-Rhys parameter, mode frequency, and correlation time, denoted as $\lambda_{j}$, $\gamma_{j}$, $\omega_{j}$, respectively. Relevant parameter values are detailed in the [10, 9, 41, 40].
The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport Green’s function can be expressed as
$\mathcal{G}_{a_{1}a_{1},a_{3}a_{3}}(t)$ $=[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
$=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (19)
where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors that satisfy $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D=\chi^{L}\chi^{R}$.
The intra-manifold exciton coherence Green’s function is given by
$\mathcal{G}_{a_{1}a_{2}}(t)$ $=\exp{[-i\omega_{a_{1}a_{2}}t-\gamma_{a_{1}a_{2}}t]}$ (20)
where we defined the pure dephasing parameter
$\tilde{\gamma}_{a_{1}a_{2}}=\gamma_{a_{1}a_{2}}-\frac{1}{2}(K_{a_{1}a_{1},a_{1}a_{1}}+K_{a_{2}a_{2},a_{2}a_{2}}).$ (21)
and the dephasing parameters $\gamma_{a_{1}a_{2}}=-K_{a_{1}a_{2},a_{1}a_{2}}$.
Figure 9: Histograms describing distribution of spectral gaps corresponding for $\omega_{eg}$, $\omega_{fg}$, and $\omega_{fe}$ at the top, middle and bottom panel respectively. Each energy bin have a width of $200\,\text{cm}^{-1}$. An increase in the number of counts indicates corresponding increase in inter-manifold transitions. For a detailed discussion, refer to Appendix B.
The state-specific line-broadening functions are defined as
$\gamma_{a_{1}}(t)$ $=\sum_{a_{2}}C(\omega_{a_{1}a_{2}})\sum_{m}T_{ma_{1}}T_{ma_{1}}T_{ma_{2}}T_{ma_{2}}$ (22)
The dissipation function, $C(\omega_{ab})$, is the Fourier transform of the phonon correlation function evaluated at exciton transition frequencies $\omega_{ab}$, expressed as
$C(\omega_{ab})=\int_{0}^{\infty}dt\exp(i\omega_{ab}t)$
$\qquad\times\int\frac{d\omega}{2\pi}J_{\text{ph}}(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
It indicates that the line broadening functions are determined by the phonon spectral function, and exciton gaps. The inter-manifold coherence Green’s functions, contains excitonic resonances of the type: $iz_{a_{1}a_{2^{\prime}}}=\omega_{a_{1}a_{2^{\prime}}}-i\gamma_{a_{1}a_{2^{\prime}}}t$, where the primed indices indicate states belonging to different manifolds. The values of the inter-manifold dephasing are estimated from the expression of the line broadening by replacing respective manifold terms with primed indices. Correlations between the source of driving and exciton-phonon dissipation was neglected [12].
Appendix B Note on the exciton energy distribution: one and two-exciton manifold
As stated in section II), the exact-diagonalization step yields $14$, one exciton states (denoted $|e_{j}\rangle$) and $105$ two-states (denoted $|f_{k}\rangle$). The distribution of transition energy gaps or spectral gaps for each of these manifolds can be evaluated using a pair of exciton energies belonging to different manifolds. For transitions between the ground and one-exciton states, ranging from $\omega_{e_{1}g_{1}}=E_{e_{1}}$ to $\omega_{e_{14}g_{1}}=E_{e_{14}}$, there are $14$ spectral gaps. The same between the ground and two-exciton states are $105$, which ranges from $\omega_{f_{1}g_{1}}=E_{f_{1}}$ to $\omega_{f_{105}g_{1}}=E_{f_{14}}$. In comparison, spectral gaps between the one and two exciton manifold spans from $\omega_{f_{1}e_{14}}=13909.6\,\text{cm}^{-1}$ to $\omega_{f_{105}e_{1}}=16847.3\,\text{cm}^{-1}$ contributing to a $14\times 105$, a polynomial large number of spectral gaps. A majority of them are dipole-allowed. These gaps appear in the expressions of exciton resonances via Green’s functions. The distribution of spectral gaps are visualized via three histograms in Fig. 9.
In the Fig. 4, the simulation considered a special situation: the pump photon energies are downconverted to degenerate entangled photon pairs. The energies of the sequentially corresponds to the each spectral gaps corresponding to each state in the the two-exciton manifold. For the chosen two examples of two exciton states $f$ and $f$, described in
Appendix C Note on the observable: two-exciton population distribution
Here, we present an outline for the derivation of the signal presented in Section III.1. The observable is defined using the projection operator onto the desired two-exciton state, averaged over the time-dependent density operator. Assuming the perturbative regime of exciton-driving field interaction to hold, the density operator can be expanded to the fourth-order in incoming field-exciton interaction Hamiltonian to obtain
$\rho_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
$=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}dt_{4}\int^{\infty}_{0}dt_{3}\int^{\infty}_{0}dt_{2}\int^{\infty}_{0}dt_{1}\langle P_{+}\mathcal{G}(t_{4})d_{R}\mathcal{G}(t_{3})d_{R}\mathcal{G}(t_{2})d_{L}^{\dagger}\mathcal{G}(t_{1})d^{\dagger}_{L}\rangle$
$\qquad\langle E^{\dagger}(t-t_{4})E^{\dagger}(t-t_{4}-t_{3})E(t-t_{4}-t_{3}-t_{2})E(t-t_{4}-t_{3}-t_{2}-t_{1})\rangle$ (24)
In the expression, the variables $\tau_{j}$ and $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ denote the exciton-photon interaction times and time-delays, respectively. The four-point correlation function of the driving field correlate the four-point correlation function of the exciton operators via the time integrals. Expanding the exciton correlation function in the basis of one and two-exciton states, carrying out the averaging over the phonons and inserting the frequency domain variables lead the expression of the signal
$\rho_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}^{*}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fe^{\prime}}(-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fg}(\omega_{2}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{eg}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{fe^{\prime\prime}}(+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}e^{\prime\prime},ee}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{fe^{\prime}}(+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{ee^{\prime}}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime\prime}}^{*}d_{fe^{\prime\prime}}d_{eg}d_{eg}^{*}\times\mathcal{G}_{ff}(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}f}(-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}e^{\prime\prime},ee}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe}^{*}d_{fe^{\prime}}d_{e^{\prime}g}d_{eg}^{*}\times\mathcal{G}_{ff}(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{ef}(-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime}e}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})\Big\}\Big\}$ (25)
The expression contains three types of exciton Green’s functions that describes inter-manifold coherence, intra-manifold coherence and transport. The first term represents the two-exciton coherence-mediated pathways, the second and fourth terms represents one-exciton transport-mediated pathways and the third and fifth terms represents the one-exciton coherence-mediated pathways. We also note that the Green’s functions contains frequencies of the driving field in their arguments. The expression can be simplified by plugging in the expressions of the exciton Green’s functions. It yields
$\rho_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{\frac{d_{fe^{\prime}}}{-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1}-z_{ff}}\frac{d_{e^{\prime}g}}{-\omega_{3}+\omega_{2}+\omega_{1}-z_{fe^{\prime}}}\frac{d_{fe}^{*}}{+\omega_{2}+\omega_{1}-z_{fg}}\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}$
$+\frac{d_{fe^{\prime\prime}}}{-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1}-z_{ff}}\frac{d_{fe^{\prime\prime}}^{*}}{\omega_{2}-\omega_{3}+\omega_{1}-z_{fe^{{}^{\prime\prime}}}}$
$d_{e^{\prime}g}\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}\frac{1}{-\omega_{3}+\omega_{1}-z_{p}}\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}\frac{1}{-\omega_{3}+\omega_{1}-z_{ee^{\prime}}}\right)\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}$
$+\frac{d_{fe^{\prime\prime}}^{*}}{+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1}-z_{ff}}\frac{d_{fe^{\prime\prime}}}{-\omega_{4}-\omega_{3}+\omega_{1}-z_{fe^{\prime\prime}}}$
$d_{e^{\prime}g}\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}\frac{1}{-\omega_{3}+\omega_{1}-z_{p}}\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}\frac{1}{-\omega_{3}+\omega_{1}-z_{ee^{\prime}}}\right)\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}\Big\}\Big\}$ (26)
A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eq. III.1. In the final step, tedious yet straightforward frequency integrations yields the expression below
$\rho_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(+z_{fe^{\prime\prime}}-z_{ff})E(-z_{p}+z_{fe^{\prime\prime}})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{eg}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
$+d_{fe^{\prime}}d_{fe^{\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(+z_{fe^{\prime}}-z_{ff})E(-z_{ee^{\prime}}+z_{fe^{\prime}})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{e^{\prime\prime}f}+z_{ff})E(z_{p}-z_{e^{\prime\prime}f})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{ee^{\prime}}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big\}\Big]$ (27)
This expression is valid for any general correlated photonic source and can be used in numerical simulation.
Appendix D Entangled photon correlation function
Here we examine the time-frequency properties entangled photon pairs. We plot the square of the joint spectral amplitude function, for a parameter regime comparable to that used in the simulation in Section III.
Figure 10: The squared joint spectral amplitudes of the entangled photon pairs are displayed, illustrating variations in the entanglement time parameter ($\tilde{T}_{\text{ent}}$) and the temporal width of the SPDC pump ($\tau_{0}$). They have noticeable implications for the nature of the frequency (anti-)correlations between the signal and idler pair, a feature that can be utilized for spectroscopic applications.
Appendix E Note on the observable: two-photon counting signal
Here, we present a brief derivation of the photon coincidence counting signal. We follow the original steps outlined in [17, 15]. The final expression is already presented in Eq. III, and appears as
The bare two-photon counting signal is expressed as
$\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\int_{-\infty}^{t_{1}^{\prime}}dt_{1}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
$\times\int_{-\infty}^{t_{2}^{\prime}}dt_{2}\int_{-\infty}^{t_{2}^{\prime}+\tau_{2}}dt_{4}\langle V_{R}^{\dagger}(t_{4})V_{R}^{\dagger}(t_{3})V_{L}(t_{1})V_{L}(t_{2})\rangle$
$\times\sum_{s,s^{\prime}}\sum_{i,i^{\prime}}\langle E_{s^{\prime}R}(t_{4})E_{i^{\prime}R}(t_{3})E_{s^{\prime}R}^{\dagger}(t_{2}^{\prime}+\tau_{2})E_{i^{\prime}R}^{\dagger}(t_{1}^{\prime}+\tau_{1})$
$\times E_{iL}(t_{1}^{\prime})E_{sL}(t_{2}^{\prime})E_{iL}^{\dagger}(t_{1})E_{sL}^{\dagger}(t_{2})\rangle.$ (28)
The field correlation term is averaged over the detector modes in vacuum. The exciton correlation function is averaged over the the two-exciton states that corresponds to the population terms of the density operator following the transport. We assumed that the leading order contributions to the photon counting signal arise from expansion of the density operator to the fourth order in emitted photon field. The emitted field modes are expanded into plane wave mode and following the standard algebraic manipulations we obtain
$\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\mathcal{D}^{2}(\omega_{1})\mathcal{D}^{2}(\omega_{2})$
$\times\langle V_{R}^{\dagger}(t_{2}^{\prime}+\tau_{2})V_{R}^{\dagger}(t_{1}^{\prime}+\tau_{1})V_{L}(t_{1}^{\prime})V_{L}(t_{2}^{\prime})\rangle.$ (29)
It highlights the fact that the four-point exciton correlation function weighted by the density of detector modes evaluated at emitted frequencies. The exciton correlation function can be expanded further to yield the Liouville space pathways depicted in the diagram in Fig. 3. The expansion also yields the expressions in  III.3.